OpenStudy (anonymous):

Write the formula given the following arithmetic sequences. tn = 1/2, 1 3/4, 3, 4 1/4, ...

OpenStudy (anonymous):

help.!

OpenStudy (radar):

It appears it is an arithmetical sequence with the difference of 5/4 , in other words the next number would be: 4 1/4 +5/4 = 22/4 =5 1/2